OpenStudy (anonymous):

Please see if I did this right?? Which of the following is a solution to the equation x - 5/7 = - 3/7. My answer is 1 1/7. Thank u!! :)

OpenStudy (anonymous):

X-5/7=-3/7 add -5/7 to both sides. x-5/7+5/7=-3/7+5/7 x=(-3+5)/7=2/7 good luck

OpenStudy (anonymous):

oh my... appreciate the explanation. thank u so much. this will help me with other problems as well. :)